An office heater puts out 2030 Btu

(British thermal units) of energy in the form of heat per hour. Given that 1 Btu=1.055 kJ,

how many megajoules of energy in the form of heat can be produced per year (365 days) by the heater, assuming continuous operation?


1
Expert's answer
2021-02-09T03:35:13-0500

1 Btu =1.055KJ

Energy in 1 hour = 2030 Btu

Energy in 1 hour = 2030 × 1.055 KJ = 2141.65KJ

1 day = 24 hours

Energy in 1 day = 2141.65 × 24 = 51399.6 KJ

1 year = 365 days

Energy in 1 year or 365 days = 51399.6 × 365 = 18760854 KJ

Now covers 18760854 KJ into megajoules

1 KJ = 0.001 megajoules

18760854 KJ = 18760.854 megajoules

In one year 18760.854 megajoules of energy in the form of heat can be produced per year (365 days) by the heater, assuming continuous operation